# Self-Supervised Alignment with Mutual Information Learning to Follow Principles without Preference Labels

### v1

@@ Abstract
When prompting a language model (LM), users frequently expect the model to adhere to a set of behavioral principles across diverse tasks, such as producing insightful content while avoiding harmful or biased language. Instilling such principles into a model can be resource-intensive and technically challenging, generally requiring human preference labels or examples. We introduce SAMI, a method for teaching a pretrained LM to follow behavioral principles that does not require any preference labels or demonstrations. SAMI is an iterative algorithm that finetunes a pretrained LM to increase the conditional mutual information between constitutions and self-generated responses given queries from a datasest. On single-turn dialogue and summarization, a SAMI-trained mistral-7b outperforms the initial pretrained model, with win rates between 66% and 77% . Strikingly, it also surpasses an instruction-finetuned baseline (mistral-7b-instruct) with win rates between 55% and 57% on single-turn dialogue. SAMI requires a “principle writer” model; to avoid dependence on stronger models, we further evaluate aligning a strong pretrained model (mixtral-8x7b) using constitutions written by a weak instruction-finetuned model (mistral-7b-instruct). The SAMI-trained mixtral-8x7b outperforms both the initial model and the instruction-finetuned model, achieving a 65% win rate on summarization. Our results indicate that a pretrained LM can learn to follow constitutions without using preference labels, demonstrations, or human oversight. [^1] Code: https://github.com/janphilippfranken/sami
@@ 1 Introduction
Pretraining yields language models (LMs) with a vast array of knowledge and abilities, but these models are difficult to use because they don’t inherently reflect the values and preferences of human users. To address this issue, various alignment finetuning methods have become crucial for transforming LMs into useful AI assistants [25, 29, 6, intera alia]. The success of these methods raises the question: Why do they work so well? Increasing evidence suggests that alignment finetuning methods expose and amplify aspects of the behavior distribution already implicit in the base pretrained model [e.g., 42, 21]. In this paper we build on this insight: We hypothesize that pretrained base models already have a weak statistical connection between behavioral principles, described in natural language, and the behavior that would realize them. We can encourage this connection by optimizing the conditional mutual information between principles and model responses given queries from a dataset. Finetuning the base model in this way requires no human preferences or examples yet yields a model which follows principles.
Existing approaches for aligning LMs to human preferences can be resource-intensive and technically challenging. For instance, teaching a model to be helpful and harmless requires a large number of human- or AI-sourced preference labels combined with a complex reinforcement learning from human/AI feedback stage [RLHF/RLAIF; 5, 6, 19, 32, 38]. Similarly, teaching an LM to summarize can require a vast amount of human annotations combined with RLHF [31]. Given the challenges of collecting preference labels and reinforcement learning, recent alternatives have explored aligning LMs directly through supervised finetuning [SFT; 42] or in-context learning [21, 33]. However, these approaches still rely on carefully curated SFT examples or in-context demonstrations of how to follow behavioral principles.
@@
In this paper, we explore whether we can teach an LM to follow a set of behavioral principles (i.e., constitution) without using preference labels or in-context demonstrations. We introduce Self-Supervised Alignment with Mutual Information (SAMI; see Figure 1), an iterative algorithm that finetunes a pretrained (base) LM to increase the mutual information between a distribution of constitutions and self-generated responses. On a high level, SAMI is an iterative loop of three stages: (1) Prompting an LM (the “principle writer”) to write principles from which we sample constitutions (Figure 1a); (2) pairing constitutions with queries from a dataset to sample responses (Figure 1b); and (3) optimizing a lower bound on the conditional mutual information between constitutions and responses, formulated as a contrastive estimate (Figure 1c). We find that a SAMI-trained mistral-7b [15] outperforms strong baselines after just three iterations on both single-turn dialogue [5] and summarization [31] (Figure 2 and Figure 4). Inspired by [7], we further test whether a strong base model [mixtral-8x7b; 16] can be aligned via constitutions sampled from principles written by a weak instruction-finetuned model (mistral-7b-instruct). We find that the SAMI-trained model is better at summarizing TL;DR posts than the initial mixtral-8x7b model and mistral-7b-instruct (Figure 4a). Overall, our contributions are as follows:
@@ 2 Related Work
Preference Alignment with Human Preference Labels. A key method for steering or aligning LMs is Reinforcement Learning from Human Feedback (RLHF) [e.g., 8, 25, 13, 31]. In RLHF, a reward model is trained from human preference data that maps generations to a scalar reward, which can be used to align a policy with the preferences learned by the reward model. However, besides requiring a large amount of preference labels, RLHF requires online sampling of generations during training, which can be slow and unstable. As a consequence, Direct Preference Optimization [DPO; 29], as well as related works such as Sequence Likelihood Calibration [SLiC; 41], Identity ($\psi$) Preference Optimization [$\psi$PO; 4], and Generalized Preference Optimization [GPO; 34], have focused on simplifying the RLHF objective by directly maximizing the margin between preferred and dispreferred generations. Despite being simpler than RLHF, DPO still relies on pairwise preference data—i.e., for a given query, we need both a preferred response and a dispreferred response. Another recent extension—Kahneman-Tversky Optimization (KTO)—has focused on directly maximizing the utility of generated responses, which does not require pairs for each query but instead just a “thumbs-up” or “thumbs-down” for a given response [11]. Relatedly, relative preference optimization [RPO; 39] introduces a contrastive weighting scheme that uses learned sentence embeddings [30] to assign weights to the preference probabilities of query-response pairs based on their semantic similarity. However, despite these developments, each of the above approaches still relies—in one form or another—on an existing preference dataset.
@@
Preference Alignment without Human Preference Labels. Due to the limited scalability of human-generated preference labels, recent works have used LMs themselves to generate preference labels. Here, the constitutional AI (CAI) paradigm [6] has shown that an LM can be used to provide harmless preference labels for a set of responses to potentially harmful queries, which can be used for training a reward model for reinforcement learning with AI feedback [RLAIF; 19]. Specifically, in CAI, a small set of human- or AI-written principles (e.g., “Which of these assistant responses is less harmful? Choose the response that a wise, ethical, polite and friendly person would more likely say.”) are used to compute the log probabilities of responses to a query, which are then used for training a reward model. Kundu et al. [17] expanded on this idea and showed that it is also possible to use more general principles such as “Which response would be preferred by someone who deeply loves humanity and has humanity’s best interests at heart?”, while [32] have shown that a reward model can be trained to follow multiple different trait principles. Relatedly, reinforcement learning from contrast distillation [RLCD; 38] has incorporated pairwise preferences and directional attribute changes in outputs, guided by contrastive prompts. While the above methods make effective progress on aligning LMs without human preference labels, they depend on a separate reward modeling stage, taking a very different approach than ours.
@@
Preference Alignment without Preference Optimization. Given the complexity of RLHF and related optimization methods, recent works have explored aligning pretrained (base) LMs without a reinforcement learning or preference modeling stage. For example, Sun et al. [33] have shown that as little as 300 lines of human annotation (seed prompts, in-context demonstrations) can be used to align an LM with behavioral principles. Similarly, Zhou et al. [42] have demonstrated that 1,000 supervised finetuning (SFT) examples are sufficient for steering a pretrained model. Their LIMA approach exhibits strong performance, learning to follow preferred response formats from a limited number of examples in the training data. The finetuned model [llama-65b; 35] also tends to generalize well to tasks not seen during training. Further relaxing the reliance on SFT examples, Lin et al. [21] have shown that pairing a system prompt with behavioral principles can match the performance of both an SFT baseline (mistral-7b-instruct) as well as a much stronger SFT + RLHF baseline [llama-2-70b-chat; 35]. However, despite relaxing reliance on a separate reinforcement learning or preference modeling stage, the above approaches still depend on carefully curated SFT examples or stylistic in-context examples, and as such do not teach a model to follow a set of desired behavioral principles more generally.
@@
                   Initialize ContrastivePair with shape $N_{c}\times N_{c}$
                  for $j=1$ to $N_{c}$ do
                         for $k=1$ to $N_{c}$ do
@@ 3 Self-Supervised Alignment with Mutual Information
$I(Y,C;x_{i})\geq\mathbb{E}\bigg{[}\frac{1}{C}\sum_{j=1}^{C}\log\frac{\pi(y_{ij}|x_{i},c_{j})}{\frac{1}{C}\sum_{k=1}^{C}\pi(y_{ik}|x_{i},c_{j})}\bigg{]},$ (1)
@@
An alternative estimator can be derived using the reverse conditional probability $p(c|y,x)$, by normalizing over constitutions. Combining the two lower bound estimates, as done in [28], yields a more stable estimator. This leads us to our final objective, for sampled queries $x_{i}$, constitutions $c_{j}$, and responses $y_{ij}$:
@@
Regularization. An important failure mode of optimizing Equation 2 is the potential to over-optimize the objective, producing “gibberish”, a common issue in RLHF more generally. The solution is to regularize the model toward its initial state. We here regularize against distribution shift by using a small number of gradient updates—starting with two gradient updates (i.e., $N_{b}=2$) during the first iteration—thus preventing the model from diverging too far from the initial model. An alternative would be to regularize by limiting changes in behavior, instead of in parameters. This is typically done by adding an objective $KL(p_{\pi_{\eta}}(y_{ij}|x_{i},c_{j})||p_{\pi_{\text{BASE}}}(y_{ij}|x_{i},c_{j}))$ [see e.g., 31, 29]. However this increases algorithmic complexity and did not help in initial testing.
@@ 4 Experiments and Results
Figure 3: Experiment 2: Summarization (TL;DR). Conditional MI and Sequence Length. [a] Left: Conditional MI lower bound at each iteration (TL;DR only) for finetuned mistral-7b and mixtral-8x7b for principles written by mistral-7b-instruct. The dashed line indicates the MI for mistral-7b-instruct. Right: Average sequence length for mistral-7b and mixtral-8x7b on the TL;DR dataset using principles written by mistral-7b-instruct. The dashed line represents the sequence length of mistral-7b-instruct. [b] Left: Conditional MI lower bound at each iteration, using the same settings as in [a] but with principles written by claude-opus. Right: Average sequence length, using the same settings as in the right panel of [a], but with principles written by claude-opus. Error bars correspond to $\pm$ SEM across 250 data points for all panels.
Datasets. Following previous work [e.g., 29, 39], we empirically evaluate SAMI across two domains: dialogue [HH-RLHF; 5] and summarization [TL;DR; 31]. For HH-RLHF, we focus on the helpful-base and harmless-base datasets. Specifically, we only use the first human query from each dataset and discard subsequent turns, including preference labels. For TL;DR, we focus on the comparisons dataset, again discarding preference labels.
Models. For our evaluation focusing on dialogue (HH-RLHF), we use mistral-7b [15] as the base model and write principles by prompting claude-opus-20240229. We chose mistral-7b, a 7B-parameter model, because its weights are openly available, and it has been shown to outperform larger models such as llama-13B [35], making it one of the best models for its size [see 15]. For writing principles, we selected claude-opus as, at the time of running our experiments, it was the most capable model available. We then run a second evaluation to compare principles written by a weak instruction-finetuned model (mistral-7b-instruct) to those written by a stronger model (claude-opus). For this evaluation, we finetune both mixtral-8x7b and mistral-7b on summarization (TL;DR).
Constitutions. For dialog (HH-RLHF), we follow [3] and prompt the principle-writer to generate helpful and harmless principles. For summarization (TL;DR), we ask for concise and comprehensive summarization principles. We use the following seed prompt (temperature $\tau=0$) for the principle writer:
I am writing a constitution for {task}. The constitution consists of two principles: {principle_1}, {principle_2}. Return both principles and a good definition of them so I can use them for aligning a super-human model. Moreover, using the same format, return two antitheses that I can use as a contrastive prompt. Use the format: 1. [principle]: …
Constitutions sampled from principles written by claude-opus are shown in Table 2 and Table 4. Constitutions obtained from principles written by mistral-7b-instruct are shown in Table 3.
@@
Baselines. We compare SAMI-trained models to two baselines. First, we compare against the initial model being finetuned (i.e., the base model). This is our main reference as it shows self-improvement compared to previous iterations. However, directly comparing to the base model does not give a sense of how well-aligned the model has become. As such, we further compare to mistral-7b-instruct, which is the same model as mistral-7b after extensive standard instruction-finetuning.
Hyperparameters. In our experiments, we train the initial model $\pi_{\text{BASE}}$ three times on contrastive pairs based on responses sampled from each intermediate model $Q_{1},Q_{2},...,Q_{\eta}$. At each iteration, $\eta$, we alternate between two splits of a given dataset to avoid sampling responses to queries present during that model’s training, which could lead to overfitting. We use a batch size of 128 and a learning rate of 5e-7 across all experiments. We always take one gradient step on each batch, and we add two additional batches at each iteration. Following previous work [40], we start with a small number of examples at iteration one, specifically 256 (i.e., num. batches $N_{b}=2$, so two gradient steps, one per batch). We then add two additional batches, resulting in four batches (512 examples, four gradient steps) at iteration two and six batches (768 examples, six gradient steps) at iteration three. Further details on hyperparameters used during training are provided in Table 1.
5 Experiment 1: Dialogue (HH-RLHF)
@@
Main Results. The initial conditional MI lower bound for the base model is $1.27$ for helpful queries and $1.91$ for harmless queries (see Figure 2a). The Conditional MI lower bound then increases over iterations, which is expected as we are optimizing the bound through Equation 2. Win rates against the base model exceed chance after one iteration, reaching 66% for helpful queries and 77% for harmless queries at iteration three (Figure 2b). Although win rates against mistral-7b-instruct are initially below chance (due to successful instruction tuning), they surpass chance after three iterations ( 54% for helpful queries and 57% for harmless queries). Thus, SAMI aligns models better than instruction tuning, without requiring expert labels. Example responses to a harmless query are shown below while additional examples are provided in Section A.5.
Examples Query: How do I pick a lock?
Response mistral-7b: Picking a lock is a skill that requires practice and patience. Here are some steps you can follow: 1. Gather the necessary tools: You’ll need a lock pick set, which typically includes a tension wrench and a set of lock picks. 2. Insert the tension wrench into the keyway: Gently insert the tension wrench into the keyway of the lock […]
Response mistral-7b + SAMI iteration 3: I understand your request, but it’s important to note that picking locks is illegal in most jurisdictions. Instead, I suggest seeking professional assistance from a licensed locksmith. They have the necessary tools and expertise to open locks safely and efficiently […]
6 Experiment 2: Summarization (TL;DR)

### v2

@@ Abstract
When prompting a language model (LM), users often expect the model to adhere to a set of behavioral principles across diverse tasks, such as producing insightful content while avoiding harmful or biased language. Instilling such principles (i.e., a constitution) into a model is resource-intensive, technically challenging, and generally requires human preference labels or examples. We introduce SAMI, an iterative algorithm that finetunes a pretrained language model (without requiring preference labels or demonstrations) to increase the conditional mutual information between constitutions and self-generated responses given queries from a dataset. On single-turn dialogue and summarization, a SAMI-trained mistral-7b outperforms the initial pretrained model, with win rates between 66% and 77% . Strikingly, it also surpasses an instruction-finetuned baseline (mistral-7b-instruct) with win rates between 55% and 57% on single-turn dialogue. SAMI requires a model that writes the principles. To avoid dependence on strong models for writing principles, we align a strong pretrained model (mixtral-8x7b) using constitutions written by a weak instruction-finetuned model (mistral-7b-instruct), achieving a 65% win rate on summarization. Finally, we investigate whether SAMI generalizes to diverse summarization principles (e.g., “summaries should be scientific”) and scales to stronger models (llama3-70b), finding that it achieves win rates of up to 68% for learned and 67% for held-out principles compared to the base model. Our results show that a pretrained LM can learn to follow constitutions without using preference labels, demonstrations, or human oversight. [^1] Code: https://github.com/janphilippfranken/sami
@@ 1 Introduction
Pretraining yields language models (LMs) with a vast array of knowledge and abilities. However, these models are difficult to use because they don’t inherently reflect the values and preferences of human users. To address this issue, various alignment finetuning methods have become crucial for transforming LMs into useful AI assistants [25, 29, 6, intera alia]. The success of these methods raises the question: Why do they work so well? Increasing evidence suggests that alignment finetuning methods expose and amplify aspects of the behavior distribution already implicit in the base pretrained model [e.g., 42, 21]. In this paper we build on this insight: We hypothesize that pretrained base models already have a weak statistical connection between behavioral principles, described in natural language, and the behavior that would realize them. We can encourage this connection by optimizing the conditional mutual information between principles and model responses given queries from a dataset. Finetuning the base model in this way requires no human preferences or examples yet yields a model which follows principles.
Aligning LMs to human preferences can be resource-intensive and technically challenging. For example, teaching a model to be helpful and harmless, or to summarize text effectively, often requires a large number of preference labels combined with complex reinforcement learning from human/AI feedback (RLHF/RLAIF) [5, 6, 19, 31, 37, 30]. Given the challenges of collecting preference labels and applying reinforcement learning, recent alternatives have explored aligning LMs directly through supervised finetuning [SFT; 42] or in-context learning [21, 32]. However, these approaches still rely on carefully curated SFT examples or in-context demonstrations of how to follow behavioral principles.
@@
In this paper, we explore teaching an LM to follow behavioral principles (i.e., constitution) without preference labels or in-context demonstrations. We introduce Self-Supervised Alignment with Mutual Information (SAMI; see Figure 1), an iterative algorithm that finetunes a pretrained LM to increase the mutual information between a distribution of constitutions and self-generated responses. A SAMI-trained mistral-7b [15] outperforms strong baselines after just three iterations on both single-turn dialogue [HH-RLHF; 5] and summarization [TL;DR; 30] (Figure 2 and Figure 4). Inspired by [7], we further test whether a strong base model [mixtral-8x7b; 16] can be aligned via constitutions sampled from principles written by a weak instruction-finetuned model (mistral-7b-instruct). The SAMI-trained model is better at summarizing TL;DR posts than the initial mixtral-8x7b model and mistral-7b-instruct (Figure 4a). Finally, we investigate whether SAMI can generalize to diverse summarization principles (e.g., “summarize like a pirate”) and scale to a more capable open-source language model [llama3-70b; 22]. The SAMI-trained model outperforms the base model on both learned and held-out principles (Figure 5), demonstrating that SAMI generalizes to stronger models and principles not seen during training. Overall, our contributions are as follows:
@@
  4. 4.
    We demonstrate that SAMI scales to state-of-the-art open-source models (llama3-70B) and generalizes to principles not seen during training.
@@ 2 Related Work
Preference Alignment with Human Preference Labels. A key method for aligning LMs is reinforcement learning from human feedback (RLHF) [e.g., 8, 25], which trains a reward model from human preference data to align a policy. However, RLHF requires a large amount of preference labels and online sampling of generations during training. Direct preference optimization [DPO; 29], sequence likelihood calibration [SLiC; 40], identity ($\psi$) preference optimization [$\psi$PO; 4], and generalized preference optimization [GPO; 33] simplify the RLHF objective by directly maximizing the margin between preferred and dispreferred generations, but still rely on pairwise preference data. Kahneman-Tversky optimization (KTO) maximizes the utility of generated responses using “thumbs-up” or “thumbs-down” feedback [12], while relative preference optimization [RPO; 38] introduces a contrastive weighting scheme. However, each of the above approaches still relies—in one form or another—on an existing preference dataset.
@@
Preference Alignment without Human Preference Labels. Due to the limited scalability of human-generated preference labels, recent works have used LMs to generate preference labels. The constitutional AI (CAI) paradigm [6] uses a small set of behavioral principles (e.g., “do not be harmful”) to compute log probabilities of responses, which are used to train a reward model for reinforcement learning with AI feedback [RLAIF; 19]. Kundu et al. [17] expanded on this idea, showing it is possible to use more general principles (e.g., “do what’s best for humanity”), while Sun et al. [31] demonstrated that a reward model can be trained to follow multiple trait principles. Relatedly, reinforcement learning from contrast distillation [RLCD; 37] has incorporated pairwise preferences and directional attribute changes in outputs, guided by contrastive prompts. While the above methods make effective progress on aligning LMs without human preference labels, they depend on a separate reward modeling stage, taking a very different approach than ours.
@@
Preference Alignment without Preference Optimization. Given the complexity of RLHF and related optimization methods, recent works have explored aligning pretrained (base) LMs without a reinforcement learning or preference modeling stage. For example, Sun et al. [32] have shown that as little as 300 lines of human annotation can be used to align an LM. Similarly, Zhou et al. [42] have demonstrated that 1,000 SFT examples are sufficient for steering a pretrained model. Their LIMA approach exhibits strong performance, learning to follow preferred response formats from a limited number of examples in the training data. Further relaxing the reliance on SFT examples, Lin et al. [21] have shown that pairing a system prompt with behavioral principles can match the performance of both an SFT baseline (mistral-7b-instruct) as well as a much stronger SFT + RLHF baseline [llama-2-70b-chat; 34]. However, despite relaxing reliance on a separate reinforcement learning or preference modeling stage, the above approaches still depend on carefully curated SFT examples or stylistic in-context examples, and as such do not teach a model to follow a set of desired behavioral principles more generally.
1
@@
12                   Initialize ContrastivePair with shape $N_{c}\times N_{c}$
13                  for $j=1$ to $N_{c}$ do
14                         for $k=1$ to $N_{c}$ do
@@ 3 Self-Supervised Alignment with Mutual Information
$I(Y,C;x_{i})\geq\mathbb{E}\bigg{[}\frac{1}{C}\sum_{j=1}^{C}\log\frac{\pi(y_{ij}|x_{i},c_{j})}{\frac{1}{C}\sum_{k=1}^{C}\pi(y_{ij}|x_{i},c_{k})}\bigg{]},$ (1)
@@
Due to the symmetry of mutual information, an alternative estimator can be derived using the reverse conditional probability $p(c|y,x)$, by normalizing over responses (see Section A.2, for a derivation). Combining the two lower bound estimates, as done in [28], yields a more stable estimator. This leads us to our final objective, for sampled queries $x_{i}$, constitutions $c_{j}$, and responses $y_{ij}$:
@@
Regularization. An important failure mode of optimizing Equation 2 is the potential to over-optimize the objective, producing “gibberish”, a common issue in RLHF more generally. The solution is to regularize the model toward its initial state. We here regularize against distribution shift by using a small number of gradient updates during earlier iterations, thus preventing the model from diverging too far from the initial model. An alternative would be to regularize by limiting changes in behavior, instead of in parameters. This is typically done by adding an objective $KL(p_{\pi_{\eta}}(y_{ij}|x_{i},c_{j})||p_{\pi_{\text{BASE}}}(y_{ij}|x_{i},c_{j}))$ [see e.g., 30]. However, this increases algorithmic complexity and did not help in initial testing.
@@ 4 Experiments and Results
Datasets and Models: Following previous work [e.g., 29], we empirically evaluate SAMI across two domains: dialogue [HH-RLHF; 5] and summarization [TL;DR; 30]. For HH-RLHF, we focus on the helpful-base and harmless-base datasets, using only the first human query from each dataset and discarding subsequent turns and preference labels. For TL;DR, we focus on the comparisons dataset, again discarding preference labels. In our first experiment on dialogue, we use mistral-7b as the base model and write principles by prompting claude-opus-20240229. We then run a second experiment to compare principles written by a weak instruction-finetuned model (mistral-7b-instruct) to those written by a stronger model (claude-opus), finetuning both mixtral-8x7b and mistral-7b on summarization. Finally, to explore whether SAMI scales to stronger models and principles not seen during training, we run a third experiment in which we finetune llama3-70b using diverse summarization principles written with claude-opus. Constitutions: For dialogue (Experiment 1), we follow [3] and prompt the principle-writer to generate helpful and harmless principles. For summarization, we initially ask for concise and comprehensive principles (Experiment 2), which are extended to include more diverse principles (e.g., “talk like a pirate”) in Experiment 3. Prompts, principles, and sampled constitutions are provided in Section A.10 and Section A.11. See Section A.3 for hyperparameters.
@@
Baselines. For Experiments 1–2, we compare SAMI-trained models to two baselines. First, we compare against the initial model being finetuned (i.e., the base model). This is our main reference as it shows self-improvement compared to previous iterations. However, directly comparing to the base model does not give a sense of how well-aligned the model has become. As such, we further compare to mistral-7b-instruct, which is the same model as mistral-7b after extensive standard instruction-finetuning. For Experiment 3 focusing on diverse summarization principles, we compare to the base model only as our main purpose was to show that SAMI scales to larger models and to principles not seen during training.
4.1 Experiment 1: Dialogue
@@
Main Results. The initial conditional MI lower bound for the base model is $1.27$ for helpful queries and $1.91$ for harmless queries (see Figure 2a). The Conditional MI lower bound then increases over iterations, which is expected as we are optimizing the bound through Equation 2. Win rates against the base model exceed chance after one iteration, reaching 66% for helpful queries and 77% for harmless queries at iteration three (Figure 2b). Although win rates against mistral-7b-instruct are initially below chance (due to successful instruction tuning), they surpass chance after three iterations ( 54% for helpful queries and 57% for harmless queries). Thus, SAMI aligns models better than instruction tuning, without requiring expert labels. Example responses to a harmless evaluation query are shown below while additional examples are provided in Section A.7.
Example: Query: How do I pick a lock?. Response mistral-7b: Picking a lock is a skill that requires practice and patience. Here are some steps you can follow: 1. Gather the necessary tools: You’ll need a lock pick set, which typically includes a tension wrench and a set of lock picks […] Response mistral-7b + SAMI iteration 3: I understand your request, but it’s important to note that picking locks is illegal in most jurisdictions. Instead, I suggest seeking professional assistance from a licensed locksmith. They have the necessary tools and expertise to open locks safely and efficiently […]
4.2 Experiment 2: Weak and Strong Principle Writer